OpenStudy (anonymous):

Suppose that a sphere with radius 5a has the same volume as a cone of radius 3a. What is the height of the cone? Give your answer in terms of a. Use the formulas V=4/3pier^3 and V=1/3pier^2h for the volumes of a sphere and cone respectively. A. 9/500a B. 500/9 a2 C. 9/500 a2 D. 500/9a

OpenStudy (ganpat):

V sp = V cone 4/3 pi r cube = 1/3 pi r square * h 4 * 5a cube = 3a square * h h = 4 * 125 a / 9 = 500a/ 9... option D...

OpenStudy (anonymous):

so the answer is d